Write common assembly language (RISC) programs to a) sum the first n elements of an array A, and b) compute r = ab for unsigned integers a and b. Each program will consist of a driver and a subprogram. The drivers will 1) read one or more values from the keyboard, 2) call the subprogram, and 3) print a result. The subprograms must not: ● store into memory, ● use registers $1 – $9, or ● make system calls
Answer #1

Solution a) Assembly Code for sum of first n element from array: DATA arraỵ DW 100 DUP (2) ;declare array maximum size is 100ca sumValue mov sum, eax isave the sum in eax register int 21H for display the sum SumValue: push ebp mex ebp, esp mov eax, 0b) Assemblv Code for r-a*b: DATA a DB DUP (?) declare a b DB DUP (?) declare b DB DUP (2) ;declare r . CODE mov ah, 4 systemCopyable Code:

a)

.DATA
array DW 100 DUP (?) ;declare array maximum size is 100
num DB DUP(?) ;declare Number
sum DB DUP(?)
.CODE
mov ah, 4 ;system call code for int read
INT 21H   
sub al, 30H   
mov num, al
mov ebx,0 ; array offset
mov ecx,[num] ;loop counter
LOOP:
mov ah, 4 ;system call code for int read
INT 21H   
sub al, 30H   
mov array[ebx], al ;save int in the array at offset ebx
addi ebx, 4 ;increment the offset by 4
dec ecx ;decrement the ecx counter
jnz LOOP ;jump to loop if not zero
push array ;push the array values
push num ;push given num
call SumValue
mov sum, eax ;save the sum in eax register
int 21H ;for display the sum
SumValue:
push ebp
mov ebp, esp
mov eax, 0 ;save value of sum
mov ebx, [ebp+4] ;load num
mov ecx, [ebp+8] ;load the array base address
loopSum:
mov edx, [ecx] ;load array values
add eax,edx ;add array element that is edx to eax
addi ecx, 4 ;increment offset by 4
dec ebx
jnz loopSum
mov esp, ebp
pop ebp
ret

b)

.DATA
a DB DUP(?) ;declare a   
b DB DUP(?) ;declare b   
r DB DUP(?) ;declare r
.CODE
mov ah, 4 ;system call code for read int a
INT 21H   
sub al, 30H   
mov a, al
mov ah, 4 ;system call code for read int b
INT 21H   
sub al, 30H   
mov b, al
push a ;push a value
push b ;push b value
call MULTIPLY ;call MULTIPLY function
mov [r], eax ;move the result r to eax
int 21H ;for display
MULTIPLY:
push ebp
mov ebp, esp
mov eax, [ebp+4] ; load b value
mov ebx, [ebp+8] ; load a value
mul eax,ebx ;for r = a* b that is eax = eax*ebx
mov esp, ebp
pop ebp
ret